# “Feeling that I was Collaborating with Them:” A 20-year Systematic Literature Review of Social Virtual Reality Leveraging Collaboration

### v2

@@ 4.2.3. Enabling Collaboration through Social VR Affordances
This sub-theme explores how the spatial layout, visual realism, and responsiveness of the virtual environment influence users’ interactions and support collaboration in Social VR. Rather than focusing on tools or input mechanisms, the reviewed articles emphasize how the environment’s visual design, structure, and adaptability influence where and how users interact. The findings suggest that effective virtual environments in Social VR balance familiar spatial cues with flexible design elements, enabling collaborative behaviors that can surpass those possible in the physical world (Mei et al., 2021; Osborne et al., 2023; Men and Bryan-Kinns, 2019; Freiwald et al., 2020; Weissker and Froehlich, 2021; Langa et al., 2022).
@@
Two articles found that environmental realism is most valued when it reinforces familiar spatial dynamics. Mei et al. (Mei et al., 2021) found that the absence of environmental sound, realistic tools, or background detail made the virtual workspace feel less functional to the users. However, the articles discussed that greater realism does not necessarily lead to a better user experience. Osborne et al. (Osborne et al., 2023) proposed a typology of Social VR environments to support users’ engagement: Skeuomorphic, environments designed as realistic office-like spaces; Experimental, environments designed as creative and open-ended spaces, and Prefab-based, which are template-driven but customizable by the users. Their findings suggested that adaptability is often more effective than fidelity. Therefore, environments in Social VR should be tailored to users’ goals and team dynamics to better support their collaborations, instead of focusing solely on visual realism.
@@
One article explains how layout and spatial configuration shape users’ access, privacy, and coordination in Social VR. In a study on collaborative music creation, Men et al. (Men and Bryan-Kinns, 2019) found that participants favored publicly visible personal spaces that balanced individual work zones with group awareness. They also revealed trade-offs in spatial arrangements: positioning avatars side-by-side offered shared perspectives but limited equal access to shared interactive virtual tools, while face-to-face setups improved visibility for all participants but required redesigned interfaces to maintain usability. These findings underscore how spatial affordances—such as visibility, proximity, and orientation—can directly impact collaborative dynamics in Social VR applications.
@@
Two articles highlighted how social VR’s affordances—such as spatial layout and 3D visibility cues—support team coordination by aligning users’ viewpoints and actions (Freiwald et al., 2020). Freiwald et al. (Freiwald et al., 2020) compared three visual indicators for representing users’ viewpoints during a collaborative task. They found that displaying a live feed of each user’s view reduced their task’s error rates, while adding a 3D cone to indicate the boundaries of each user’s viewpoint enhanced their social presence. Similarly, Weissker et al. (Weissker and Froehlich, 2021) developed a group navigation technique that allowed collocated users to teleport together while preserving their relative positions. This feature helped teams stay spatially coordinated as they explored larger virtual environments (Weissker and Froehlich, 2021).
@@
Lastly, one study addressed the technical setup required to scale larger groups and collaborative spaces in Social VR. Langa et al. (Langa et al., 2022) developed a Social VR application for multiparty volumetric meetings, focusing on accessibility and stability for larger groups. While their evaluation involved groups of two and four, the system was designed to scale beyond small teams without sacrificing responsiveness or presence. Their use of volumetric user representations—captured and rendered in 3D—enabled more lifelike, embodied interactions than traditional avatars, offering a promising direction for realistic and scalable Social VR collaboration (Langa et al., 2022).
@@
In summary, these articles suggest that the most effective Social VR applications do not simply mimic the real world. Instead, they introduce new affordances that are not possible to afford in face-to-face meetings—such as view-sharing, spatial partitioning, and adaptable layouts—to support effective collaboration. Employing these immersive affordances to augment team members’ abilities reflects a core principle in CSCW: designing systems for beyond being there by leveraging the unique possibilities of virtual environments rather than constraining them to real-world metaphors.
@@
Articles in this sub-theme examine how virtual tools and manipulable objects enable collaboration in Social VR. Unlike traditional online collaboration platforms that rely on screen-based interactions, Social VR allow users to engage with sharedtools using hand movements, gaze direction, and spatial movement, making teamwork feel more tangible and embodied. Across the reviewed articles, tools such as pens, whiteboards, 3D models, and laser pointers were not only instrumental to complete tasks but also served a social function, enabling users to coordinate actions, maintain mutual awareness, and collaboratively build shared outcomes (García et al., 2008; Dey et al., 2024; Osborne et al., 2023).
@@
Two articles examined closely coupled collaborative tasks that required users to coordinate in real-time by synchronizing and sequentially manipulating shared objects (Roberts et al., 2005; García et al., 2008). Roberts et al. (Roberts et al., 2005) designed a Gazebo-building task in which participants had to pass, plan, and co-assemble components in real-time. Garcia et al. (García et al., 2008) addressed the technical challenges in network consistency and inconsistent feedback synchronization across users in VR by introducing visual cues—such as arrows and fall indicators—that helped participants anticipate object behavior and reduced errors during their interactions.
@@
Other articles emphasized the importance of tool accessibility and ease of use for effective collaboration in Social VR. While certain features in VR (e.g., sticky notes, drawing tools) supported brainstorming and shared design tasks, platform-specific constraints (e.g., restricted permissions, complex setup processes) often hindered the collaboration experience (Osborne et al., 2023). Osborne et al. (Osborne et al., 2023) noted that AltspaceVR provided relatively few built-in tools for groups and frequently relied on external software, which sometimes disrupted users’ interaction flow. Complementing these results, Dey et al. (Dey et al., 2024) underscored how direct interaction with virtual objects in Social VR enhances immersion, making the workspace feel more tangible. Their findings suggest that physical engagement with virtual tools not only aids task performance but also strengthens users’ sense of connection to the team and the task (Dey et al., 2024).
@@
Taken together, these articles demonstrate that the design and usability of virtual tools directly affect the collaborative dynamics in Social VR. Intuitive and accessible tools can support effective teamwork by facilitating smoother interactions, idea exchange, and action coordination in coupled tasks. In contrast, complex, restrictive, or poorly designed tools introduce friction to their work and user experiences in VR. Applications that allow multiple users to jointly control, annotate, or manipulate virtual objects foster more balanced participation and fluid idea exchange. Conversely, Social VR applications with limited multi-user access or rigid fixed interaction patterns can reinforce bottlenecks or discourage contribution.
Enhancing Immersion through Spatial Audio and Movement.
@@
The last sub-theme examines how audio and movement technologies in Social VR contribute to a deeper sense of immersion and facilitate user coordination. Affordances such as spatial audio and group locomotion help users feel oriented, responsive, and connected within the shared virtual environment, supporting both awareness and collaboration (Weissker et al., 2019; Immohr et al., 2024; Ríos et al., 2018).
@@
Two articles demonstrated that spatial audio helps anchor users in the virtual environment by simulating the direction and proximity of voices and ambient sounds. Weissker et al. (Weissker et al., 2019) found that using binaural spatial audio—which simulates 3D sound perception by delivering slightly different signals to each ear—improved triadic collaboration by enhancing social presence, turn-taking behavior, and attentional alignment. Participants reported being more aware of their teammates’ positions and smoother conversational flow compared to non-spatial audio conditions (Weissker et al., 2019). However, Immohr et al. (Immohr et al., 2023) found that spatial audio did not significantly improve performance in a negotiation task, suggesting that the effectiveness of audio design depends on the type of task and the dynamics of user interaction. The processing and presentation of spatial audio also influence how immersive and credible virtual interactions feel. Immohr et al. (Immohr et al., 2024) evaluated different binaural rendering methods and found that more realistic spatial cues improved both the perceived authenticity of communication and the naturalness of user responses. These subtle auditory refinements encouraged more lifelike exchanges, particularly in tasks where directional cues and conversational nuance played a critical role.
@@
The reviewed articles introduced a range of group navigation techniques to address the challenges of coordinating movement in collaborative VR applications (Weissker, 2020; Naef et al., 2004; Weissker et al., 2019). For instance, Weissker et al. (Weissker, 2020) developed a group teleportation system that preserved user formations during their jumps. Building on this idea, Naef et al. (Naef et al., 2004) developed the “blue-c API,” which enabled synchronized group movement, allowing users to navigate the virtual environment together without losing cohesion. Weissker et al. (Weissker et al., 2019) developed “Multi-Ray Jumping,” a system designed for collocated groups that offered directional previews, seamless teleportation, and enhanced mutual awareness during shared navigation. These articles show that group movement in VR is not only about getting from one place to another but also about maintaining spatial coordination and shared context to support effective teamwork (Naef et al., 2004; Weissker et al., 2019).
@@ 5. Discussion
This review explores two decades of research on Social VR as a medium for supporting collaboration and teamwork. Through a systematic analysis, we characterized how researchers have designed, evaluated, and applied Social VR technologies to enhance remote collaboration. Using thematic analysis, we identified three overarching themes that capture how the literature has conceptualized and evaluated collaboration in social VR applications (see Table 3).
In the following subsections, we discuss the key findings from this review, highlighting how Social VR enables new forms of presence, engagement, and coordination. We reflect on how these findings extend, complement, or challenge prior research in remote collaboration—particularly computer-mediated communication (CMC) and early VR applications—and outline how Social VR introduces new questions and tensions that require more perspectives in CSCW.
@@
Am I Really There? Examines users’ sense of presence as a foundation for immersive collaboration Presence increases with natural interaction and repeated use. Physical engagement with tasks enhances immersion. Familiarity over time deepens presence.
How Am I Perceived Here? Explores the impact of avatar design on collaborative dynamics Avatar customization supports engagement and identity. Balance needed between realism and avoiding the uncanny valley. Expressive cues shape confidence and inclusion.
@@
Are We Really Together? Focuses on how social presence enhances team collaboration Full-body avatars and non-verbal cues increase connectedness. Spatial and emotional alignment improve collaboration. Social presence fosters mutual responsiveness.
Is This Collaboration Real? Addresses the perception of authenticity in virtual teamwork Natural behaviors and responsiveness feel more authentic. Collaboration quality tied to social authenticity. Realism must support, not hinder, collaborative flow.
Should I Trust This Person? Examines trust-building in virtual collaboration Only two studies explored trust directly. Trust built through social cues and personal space. Informal interactions and mimicry enhance trust.
Are We Fully Engaged? Studies the importance of social engagement Non-verbal behaviors support engagement. Shared storytelling and personal content deepen connection. Physiological synchrony linked to team cohesion.
@@
Creating Realistic and Functional Virtual Environments Explores the design of functional virtual spaces Environments should go beyond physical replication. Layouts shape access, privacy, and coordination. Customization enhances collaboration effectiveness.
Facilitating Interaction through Virtual Tools and Objects. Examines the object manipulation and tool use Tools support both task performance and coordination. Accessibility and ease of use are critical for collaboration. New tools should support shared, dynamic interaction.
Enhancing Immersion through Spatial Audio and Movement Focuses on spatial audio and movement Spatial audio improves presence and turn-taking. Coordinated movement helps group alignment. Immersive features must match the task context.
@@ 5.1. RQ1: Diverse Research from Design to Evaluation in Social VR
The reviewed corpus reveals a growing interest in understanding how Social VR can enhance collaboration. Building on foundational CSCW research that emphasized the importance of conceptualizing distance (Olson and Olson, 2000) and place (Harrison and Dourish, 1996) in virtual communication, the articles examined Social VR as a new medium for group interaction. These articles depart from other major strands of VR research focused on gaming or education. Over the two past decades, advances in hardware, software, and networking have enabled more robust multi-user experiences and triggered new opportunities to create more Social VR experiences. Notably, the rise in research on collaborative Social VR since 2022—influenced by the COVID-19 pandemic and the surge of interest in metaverse technologies—has spurred new systems and theoretical conversations about enabling teamwork in immersive 3D environments.
@@
Our systematic review also highlights how foundational constructs in VR—such as presence, social presence, and the uncanny valley—remain central to evaluating collaboration in Social VR. For example, while the uncanny valley is a known challenge in VR, the reviewed articles show how subtle imperfections in avatar realism can disrupt trust, make users feel socially uneasy, and even affect team cohesion, issues that matter less in entertainment contexts. While these concepts have traditionally been used in HCI and communication studies to assess user experience, their application is expanding to multi-user contexts. As VR becomes more embedded in collaboration, researchers use these frameworks to understand group perception, interaction, and relationships. Bridging these constructs with perspectives from team science (Mathieu et al., 2017) and groups in CSCW (Harris et al., 2019) can offer new theoretical pathways for understanding immersive social dynamics.
One of the key insights is that every design component in Social VR can significantly influence collaboration. Unlike traditional screen-based platforms, Social VR requires full embodied engagement and attachment to the simulated experience. Its immersive nature means that even seemingly minor design decisions—such as avatar posture, hand alignment, spatial audio calibration, or tool responsiveness—can affect users’ comfort, perception, and collaborative performance. While well-designed affordances help users feel grounded and engaged in the virtual workspace, poorly implemented features can lead to disorientation, misalignment, and disengagement.
These challenges are amplified by the immersive nature of VR. While users in video conferencing tools like Zoom can multitask, momentarily disengage, or rearrange their screens with minimal disruption, in Social VR pausing or stepping away from a group session often requires removing the headset or exiting the application, causing a complete break in presence and team cohesion. This absence of pause and exit mechanisms highlights how immersive systems create new forms of “collaborative friction” that do not exist in other communication environments (Shivener and Caravella, ). These “frictions” make Social VR applications more likely to usability breakdowns, reticence, and user frustration. As our review shows, the collaborative experience is shaped by an interplay of multiple factors at the individual, team, and environmental level, including the user’s physical setup, device usability, and familiarity with immersive technologies. Even subtle issues, like delayed hand tracking or disorienting locomotion, can lead to breakdowns in trust, participation, or group cohesion. All of these factors are tightly coupled with the application’s design choices.
Unlike web-based communication platforms—which often do not include the devices as part of their design—building Social VR applications spans multiple technical layers, from HMD ergonomics to affordance design for social interaction. Not having these different layers aligned and appropriately designed can lead to more user frustration. For example, a confusing teleportation system could hinder team navigation, reduce situational awareness, and create participation asymmetries. Earlier collaboration systems often treated social dynamics and interface design as separate concerns, but the articles show that these factors are deeply intertwined. As such, supporting effective collaboration in Social VR requires an integrated approach that considers both technological and social design elements. Developers and designers must ensure that the entire system works cohesively to support seamless and equitable collaboration.

### v3

@@ 4.2.3. Enabling Collaboration through Social VR Affordances
This sub-theme explores how the spatial layout, visual realism, and responsiveness of the virtual environment impact users’ interactions and support collaboration in Social VR. Rather than focusing on tools or input mechanisms, the reviewed articles emphasize how the environment’s visual design, structure, and adaptability influence where and how users interact. The findings suggest that effective virtual environments in Social VR balance familiar spatial cues with flexible design elements, enabling collaborative behaviors that can surpass those possible in the physical world (Mei et al., 2021; Osborne et al., 2023; Men and Bryan-Kinns, 2019; Freiwald et al., 2020; Weissker and Froehlich, 2021; Langa et al., 2022; Freeman et al., 2022; Saffo et al., 2021; Zhang et al., 2023).
@@
Two articles found that environmental realism is most valued when it reinforces familiar spatial dynamics (Mei et al., 2021; Osborne et al., 2023). Mei et al. (Mei et al., 2021) found that the absence of environmental sound, realistic tools, or background detail made the virtual workspace feel less functional to users. However, the articles discussed that greater realism does not necessarily lead to a better user experience. Osborne et al. (Osborne et al., 2023) proposed a typology of Social VR environments to support users’ engagement: Skeuomorphic, environments designed as realistic office-like spaces; Experimental, environments designed as creative and open-ended spaces, and Prefab-based, which are template-driven but customizable by the users. Their findings suggest that adaptability is often more effective than fidelity. Therefore, applications in Social VR should be tailored to users’ goals and team dynamics to better support their collaborations, instead of focusing solely on visual realism.
@@
Research explains how layout and spatial configuration shape users’ access, privacy, and coordination in Social VR (Freeman et al., 2022; Saffo et al., 2021; Men and Bryan-Kinns, 2019; Zhang et al., 2023). In a study on collaborative music creation, Men et al. (Men and Bryan-Kinns, 2019) found that participants favored publicly visible personal spaces that balanced individual work zones with group awareness. They also revealed trade-offs in spatial arrangements: positioning avatars side-by-side offered shared perspectives but limited equal access to shared interactive virtual tools, while face-to-face setups improved visibility for all participants but required redesigned interfaces to maintain usability. In sum, spatial affordances—such as visibility, proximity, and orientation—can directly impact collaborative dynamics in Social VR applications.
@@
Two articles highlighted how Social VR’s affordances—such as spatial layout and 3D visibility cues—support team coordination by aligning users’ viewpoints and actions (Freiwald et al., 2020). Freiwald et al. (Freiwald et al., 2020) compared three visual indicators to represent users’ viewpoints during a collaborative task. They found that displaying a live feed of each user’s view reduced their task’s error rates, while adding a 3D cone to indicate the boundaries of each user’s viewpoint enhanced their social presence. Similarly, Weissker et al. (Weissker and Froehlich, 2021) developed a group navigation technique that allowed collocated users to teleport together while preserving their relative positions. This feature helped teams stay spatially coordinated as they explored larger virtual environments (Weissker and Froehlich, 2021).
@@
Lastly, one study addressed the technical setup required to scale larger groups and collaborative spaces in Social VR. Langa et al. (Langa et al., 2022) developed a VR application for multiparty volumetric meetings, focusing on accessibility and stability for larger groups. While their evaluation involved groups of two and four, the system was designed to scale beyond small teams without sacrificing responsiveness or presence. Their use of volumetric user representations—captured and rendered in 3D—enabled more lifelike, embodied interactions than traditional avatars, offering a medium for realistic and scalable Social VR collaboration (Langa et al., 2022).
@@
In summary, these articles suggest that the most effective Social VR applications do not simply mimic the real world. Instead, they introduce new affordances that are not possible to afford in face-to-face meetings—such as view-sharing, spatial partitioning, and adaptable layouts—to support effective collaboration.
@@
Articles in this sub-theme examine how virtual tools and manipulable objects enable collaboration in Social VR. Unlike traditional online collaboration platforms that rely on screen-based interactions, Social VR allows users to engage with shared tools using hand movements, gaze direction, and spatial movement, making teamwork feel more tangible and embodied. Across the reviewed articles, tools such as pens, whiteboards, 3D models, and laser pointers were not only instrumental to complete tasks but also served a social function, enabling users to coordinate actions, maintain mutual awareness, and collaboratively build shared outcomes (García et al., 2008, 2019; Dey et al., 2024; Osborne et al., 2023; dos Santos et al., 2011; Nguyen et al., 2017; Rasch et al., 2024).
@@
Previous research has closely examined coupled collaborative tasks that required users to coordinate in real-time by synchronizing and sequentially manipulating shared objects (Otto et al., 2006; Roberts et al., 2005, 2004; García et al., 2008). Roberts et al. (Roberts et al., 2005) designed a Gazebo-building task in which participants had to pass, plan, and co-assemble components in real-time. Garcia et al. (García et al., 2008) addressed the technical challenges of network consistency and inconsistent feedback synchronization across users in VR by introducing visual cues—such as arrows and fall indicators—that helped participants anticipate object behavior and reduce errors during their interactions.
@@
Other articles emphasized the importance of tool accessibility and ease of use for effective collaboration in Social VR (Rasch et al., 2024; Osborne et al., 2023; dos Santos et al., 2011). While certain features in VR (e.g., sticky notes, drawing tools) supported brainstorming and shared design tasks, platform-specific constraints (e.g., restricted permissions, complex setup processes) often hindered the collaboration experience (Osborne et al., 2023; dos Santos et al., 2011; Nguyen et al., 2017; Argelaguet et al., 2011). Osborne et al. (Osborne et al., 2023) noted that AltspaceVR provided relatively few built-in tools for groups and frequently relied on external software, which sometimes disrupted users’ interaction flow. Complementing these results, Dey et al. (Dey et al., 2024) underscored how direct interaction with virtual objects in Social VR enhances the sense of presence, making the workspace feel more tangible. Their findings suggest that physical engagement with virtual tools not only aids task performance but also strengthens users’ sense of connection to the team and the task (Dey et al., 2024).
@@
Taken together, these articles demonstrate that the design and usability of virtual tools directly affect the collaborative dynamics in Social VR. Intuitive and accessible tools can support effective teamwork by facilitating smoother interactions, idea exchange, and action coordination in coupled tasks. Applications that enable multiple users to jointly control, annotate, or manipulate virtual objects promote more balanced participation and facilitate fluid exchange of ideas.
Supporting Presence through Spatial Audio and Movement.
@@
The last sub-theme examines how audio and movement features in Social VR contribute to a deeper sense of presence and facilitate user coordination. Affordances such as spatial audio and group locomotion help users feel oriented, responsive, and connected within the shared virtual environment, supporting both awareness and collaboration (Weissker et al., 2019; Immohr et al., 2024b, a; Ríos et al., 2018).
@@
The identified articles demonstrated that spatial audio helps anchor users in the virtual environment by simulating the direction and proximity of voices and ambient sounds (Deacon and Barthet, 2023; Weissker et al., 2019). Weissker et al. (Weissker et al., 2019) found that using binaural spatial audio—which simulates 3D sound perception by delivering slightly different signals to each ear—improved triadic collaboration by enhancing social presence, turn-taking behavior, and attentional alignment. Participants reported being more aware of their partner’s position and smoother conversational flow compared to non-spatial audio conditions (Weissker et al., 2019). However, Immohr et al. (Immohr et al., 2023) found that spatial audio did not significantly improve performance in a negotiation task, suggesting that the effectiveness of audio design depends on the type of task and the dynamics of user interaction. The processing and presentation of spatial audio also influence users’ sense of presence and the credibility of virtual interactions. Immohr et al. (Immohr et al., 2024b) evaluated different binaural rendering methods and found that more realistic spatial cues improved both the perceived authenticity of communication and the naturalness of user responses. These subtle auditory refinements encouraged more lifelike exchanges, particularly in tasks where directional cues and conversational nuance were critical.
@@
The reviewed articles introduced a range of group navigation techniques to address the challenges of coordinating movement in collaborative VR applications (Weissker, 2020; Naef et al., 2004; Weissker et al., 2019; Rasch et al., 2023; Sarmiento et al., 2014). For instance, Weissker et al. (Weissker, 2020) developed a group teleportation system that preserved user formations during their jumps. Building on this idea, Naef et al. (Naef et al., 2004) developed the “blue-c API,” which enabled synchronized group movement, allowing users to navigate the virtual environment together without losing cohesion. Weissker et al. (Weissker et al., 2019) developed “Multi-Ray Jumping,” a system designed for collocated groups that offered directional previews, seamless teleportation, and enhanced mutual awareness during shared navigation. These articles show that group movement in VR is not only about getting from one place to another but also about maintaining spatial coordination and shared context to support effective teamwork (Naef et al., 2004; Weissker et al., 2019).
@@ 5. Discussion
This review synthesizes two decades of work on Social VR for collaboration. We identified three overarching conceptual levels that capture how prior work conceptualized collaboration in Social VR (see Table 3): individual experience (i.e., presence, identity/avatars, comfort), team dynamics (i.e., social presence, collaboration realness, trust, engagement), and affordances (i.e., space/layout, shared tools/objects, audio and movement). Although patterns emerged, the findings were mixed: some studies reported stronger feelings of “being there” and smoother coordination, while others noted flat presence scores, uncanny or awkward avatar behavior, and task-dependent effects of audio and movement. These results underscore the need to examine when Social VR helps or falls short.
In the following subsections, we discuss the key findings of this review with two goals: to identify the intersections and open debates that emerge from these identified studies, and to critically discuss the potential benefits and limitations of leveraging Social VR to support effective collaboration.
@@
Am I Really There? Examines users’ sense of presence as a basis for collaboration Presence improves with intuitive action and natural expression. VR adaptation alone did not raise presence in some studies. Missing nonverbal cues and tool friction reduce feeling ’there’.
How Am I Perceived Here? Explores how avatar design shapes collaboration Customization supports engagement and identity exploration. Near-human realism can trigger discomfort (uncanny valley). Mismatch between self-presentation and perceived image lowers confidence.
@@
Are We Really Together? How social presence enhances teamwork Shared gaze/gestures/spatial audio aid turn-taking and mutual awareness. Emotional connection rises with visible, timely nonverbal cues. Weak social presence makes teamwork feel like a phone call.
Is This Collaboration Real? Perception of authenticity in virtual teamwork Behavioral realism (timely gaze, responsive motion) matters more than looks. Transformed Social Interaction can increase cohesion when cues are aligned. Realistic replication can boost authenticity but also introduce usability challenges.
Should I Trust This Person? Trust-building in Social VR Trust has received little attention in Social VR research, with only two direct studies to date. Trust built through social cues and personal space. Informal interactions (e.g., break areas) help bonding.
Are We Fully Engaged? Social engagement during collaboration Augmented avatar cues can raise perceived engagement. Sharing personal content or self-disclosure deepens connection. Physiological synchrony (e.g., heart rate) shows deeper engagement.
@@
Creating Realistic and Functional Virtual Environments Design of spaces for work together Adaptability matters more than realism; go beyond replication. View-sharing and room layouts help balance private work with group coordination. Customization enhances collaboration effectiveness.
Facilitating Interaction through Virtual Tools and Objects. Object manipulation and tool use Tools support both task work and coordination. Multi-user authoring, annotation, and export/persistence matter. Platform constraints and setup complexity break flow.
Supporting Presence through Spatial Audio and Movement Role of spatial audio and locomotion Spatial audio improves awareness and turn-taking, but task effects vary. Group teleportation keeps teammates aligned. Movement cues can align behavior even when presence scores stay flat.
Table summarizing three overarching themes and associated sub-themes identified in the literature on Social Virtual Reality collaboration. For each sub-theme, the table lists a brief description of the focus area and a set of key insights reported across prior studies, organized under the categories of individual experience, team dynamics, and collaboration affordances.
@@ 5.1. RQ1: Diverse Research from Design to Evaluation in Social VR
Regarding how previous studies have addressed team collaboration in Social VR (RQ1), our corpus reveals the potential of Social VR as a distinct medium enabling embodied group interactions unfolding online. These articles diverge from other major strands of VR research, which focus on gaming or education, and emphasize the importance of fidelity for collaboration at the individual, team, and environmental level. The surge in research on collaborative Social VR since 2022— the COVID-19 pandemic and the renewed interest in ‘metaverse’ technologies—has spurred both novel systems and theoretical conversations about teamwork in immersive 3D environments.
@@
Our review highlights how core constructs in VR—such as presence, social presence, and the uncanny valley—shape collaboration in Social VR. For example, while the uncanny valley is a known VR challenge, the reviewed articles show how subtle imperfections in avatar realism can disrupt trust, make users feel socially uneasy, and reduce team cohesion, less evident in entertainment contexts. While these concepts were traditionally used in HCI and communication studies to assess user experience, their use is expanding to multi-user contexts. Bridging these constructs with perspectives from team science (Mathieu et al., 2017) and groups in CSCW (Harris et al., 2019) can offer new ways to understand social dynamics in immersive environments.
When assessing what supports collaboration in Social VR, these studies converge on a mechanism that we term interaction fidelity. Unlike gaming or training, which can afford lower fidelity, collaboration in Social VR depends on higher realism in avatars and environments, approaching in-person expectations (Ball, 2022). The teams from these studies benefited when Social VR allowed them to arrange information in space, share viewpoints, and manipulate common objects with predictable, low-latency responses. They struggled when hand alignment, locomotion, or tools were inconsistent. Avatar design showed a similar trend: expressive and legible behaviors enabled natural interaction, whereas near-human appearances without matching behavior trigger discomfort. Affordances were also critical. Audio and movement features were most useful when they clarified members’ attention and turn-taking, though effects were task-dependent. For example, spatial audio improved coordination in group discussions, yet offered little benefit in negotiation tasks. Overall, these results suggest that Social VR is more effective for tasks that hinge on spatial arrangement and embodied interaction—such as manipulating data objects (Olaosebikan et al., 2022) or co-designing artifacts (Mei et al., 2021)—while providing fewer advantages for tasks where such affordances matter less, such as negotiation or conversation tasks (Immohr et al., 2023).